OpenStudy (anonymous):

Write an equation in the point-slope form of the line through point j(4,1) with slope -4 ? How do i do this?:O

OpenStudy (anonymous):

start with \[y-1=-4(x-4)\] and solve for \(y\)

OpenStudy (anonymous):

Thanks, how do i solve for y?:)

OpenStudy (anonymous):

1) distribute on the right 2) add \(1\) to both sides

OpenStudy (anonymous):

alright so then i would get y= -3 (x - 4) right?:)